Actinium-225 has a half-life of 10 days. Suppose we have an initial amount of 100 grams of actinium-225. How much would be present after 30 days
Answer by stanbon(75887)

Actinium-225 has a half-life of 10 days. Suppose we have an initial amount of 100 grams of actinium-225. How much would be present after 30 days
-----
A(t) = Ao*(1/2)^(t/10)
A(10) = 100*(1/2)^(30/10)
A(10) = 100*(1/2)^3 = 12.5 grams
--------
Cheers,
Stan H.
------------